OpenStudy (anonymous):

help. Input the equation of the given line in slope-intercept form. The line with m = 3 and b = -2

OpenStudy (anonymous):

The slope intercept form of a line is given as y=mx+b substituting m = 3 and b = -2 in y=mx+b we find y=3x+(-2) y=3x-2 which is the required eq of the line.